how long does it take for a deposit of $900 to double at 2% compounded continuously
Answer by lwsshak3(11628)

how long does it take for a deposit of $900 to double at 2% compounded continuously
formula for continuous compounding: A=Pe^rt, P=initial investment, r=interest rate, t=number of years, A=amt after t-years.
For given problem:
A/P=e^rt=2
rt*lne=ln2
lne=1
t=ln2/r=ln2/.02=34.66
how long does it take for a deposit of $900 to double? about 35 yrs